Circular No. 8009

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hm IN NGC 4016 T. Boles, Coddenham, England, reports the discovery of an apparent

supernova (mag 16.9) on an unfiltered CCD image taken on Nov. 5.158 UT in the course of the U.K. Nova/Supernova Patrol. The new object is located at α = 11^{h}58^{m}29^{s}.58, δ = +27^{o}31′45′′.1 (equinox 2000.0), which is ≈ 6′′.8 east and 1′′.0 north of the center of NGC 4016. SN 2002hm is not present on Boles’ images from Apr. 4 (limiting mag 18.5), 19, or June 1, and it is not present on Palomar Sky Survey red and blue plates.

G. Pignata and F. Patat, European Southern Observatory; and S. Benetti, Istituto Nazionale di Astrofisica, Padua, on behalf of the European RTN (cf. IAUC 7961, 7987), report that full reduction of a spectrum (range 380–880 nm; resolution 1.1 nm) of SN 2002hm, obtained at Calar Alto on Nov. 6.19 UT with the 2.2-m reflector (+ CAFOS), shows it to be a type-II supernova similar to SN 1992H (Clocchiatti et al. 1996, A.J. 111, 1286), ∼ 1 month past maximum. The spectrum is dominated by hydrogen lines of the Balmer series, Ca II, and Fe II, accompanied by well-developed P-Cyg profiles, overimposed on a rather blue continuum. The photospheric expansion velocities deduced from the minima of absorption throughs of Hα and Hβ are 9400 and 8000 km/s, respectively. The recession velocity at the supernova position, deduced from the narrow Hα line from the underlying H II region, is 3560 km/s.

SUPERNOVA 2002hn IN NGC 2532 D. Hutchings and W. Li, University of California, Berkeley (UCB),

report the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova (mag ∼ 16.8) on unfiltered KAIT images taken on Nov. 5.5 and 6.5 UT. SN 2002hn is located at α = 8^{h}10^{m}14^{s}.95, δ = +33^{o}57′19′′.4 (equinox 2000.0), which is 0′′.5 west and 4′′.6 south of the nucleus of NGC 2532. A KAIT image taken on Oct. 30.5 showed the new object at mag 17.3, while an image on Oct. 21.5 showed nothing at this position (limiting mag ∼ 19.0). SN 1999gb also occurred in NGC 2532. R. Foley, Li, and A. V. Filippenko, UCB, report that inspection of a CCD spectrum (range 310–1000 nm), obtained on Nov. 6 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2002hn is of type Ic, within one week of maximum brightness.

2002 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green